# Measuring Chain of Thought Faithfulness by Unlearning Reasoning Steps

### v2

@@ 6.1 Effectiveness of Unlearning
We report the results of unlearning when using the best hyperparameters for each method and dataset in Table 1. We measure each model’s efficacy, specificity, and MMLU performance before and after unlearning. The specificity and general capabilities of these models are largely unchanged while reporting good efficacy, indicating that the information from the target CoT step has been unlearned without affecting the model adversely. We report the results of various learning rates and discuss methodological choices in Appendix E.
@@ 6.2 Does Unlearning Change Predictions?
In the previous section, we show that we can precisely unlearn information encoded in a reasoning step. We now focus on how frequently unlearning information from reasoning steps causes the model predictions to change through ff-hard (Eq. 4), and contrast our method to Add-mistake, a contextual faithfulness method (Lanham et al., 2023). Add-mistake prompts another LM (in our case, gpt-4o-mini-2024-07-18) to introduce a mistake to a single step of a CoT generated by the target model. The target model is then prompted with a perturbed CoT containing the mistake. If the prediction of the model changes, the CoT is considered faithful. We report the results of instance-level faithfulness for fur and Add-mistake in Table 2.888We explore and comment other baselines in Appendix B.
@@
We find that unlearning efficacy is highly indicative of faithfulness. The Pearson correlation between average efficacy and ff-hard is high: $0.889$ with $p<0.0001$. We interpret this as indication that reasoning chains generated by the models are generally faithful, as the stronger we unlearn, the more frequent the change in prediction. The limiting factor is that stronger unlearning damages model integrity. Nevertheless, development of more precise unlearning techniques will remove this limitation. We discuss this further, along with step-level faithfulness Appendix H.
@@ 6.3 Does Unlearning Change Reasoning?
Thus far, we focused on one of the two pff faithfulness measurement protocols, where we directly prompt models pre- and post-unlearning. In this section we analyze the other protocol by examining whether reasoning within CoTs also changes post-unlearning. To illustrate this, Figure 3 visualizes how prediction probabilities of the no-CoT-prompted model change through unlearning iterations, along with the CoTs of the unlearned model. ‘Base’ refers to the model pre-unlearning. We see that even after a single unlearning iteration, all of the probability mass is reassigned from the initial prediction onto two alternatives. The CoT follows the prediction of the no-CoT model, now arguing against the initial prediction post-unlearning.
To quantitatively assess how frequently the verbalized reasoning of the model changes post-unlearning, we employ an LLM-as-a-judge (Zheng et al., 2023a) to verify if unlearning caused the generated CoT to support a different answer, indicating deeper unlearning, or if the change in model prediction is not reflected in reasoning Cohen et al. (2024). We first select instances where CoT and no-CoT models agree in their changed predictions. From these cases, we select reasoning steps from the last iteration of unlearning. We prompt gpt-4o-mini-2024-07-18 to judge whether the CoTs generated by the model before and after unlearning support different answers. We report results in Table 3 and detail our setup in Appendix G.
Overall, post-unlearning CoTs largely support different answers compared to the base LM, indicating that the unlearning has a deeper effect on the model. We believe these results further confirm the validity of our approach. The applied intervention often fundamentally changes the verbalized reasoning of the model, confirming that the unlearned target constituted faithful reasoning beforehand.
Figure 4: Heatmap produced by unlearning reasoning steps. $\Delta p$ indicates change in initial answer probability. Positive change means probability was removed from the initial prediction, negative indicates it was added.
@@ 6.4 Quantifying Step Level Faithfulness
In this section, we showcase how ff-soft (Eq. 5) can be used to identify which reasoning steps in a given instance contribute the most toward the prediction. In Figure 4 we plot heatmaps for each reasoning step, which indicate how much probability mass has been shifted to (red) or from (green) the models’ initial answer when that step was unlearned. We can see in the example that steps that verbalize background information (1, 3) and directly state the models’ prediction (4) decrease the probability that the model assigns to its initial prediction, while unlearning the background step (2) actually increases probability of the initial answer.
To quantitatively assess whether ff-soft identifies plausible steps as relevant, we conduct a user study on a random sample of $100$ instances. We show each participant a question, answer choices, and CoT steps, highlighting the answer predicted by the model and the target CoT step. We prompt the participants to annotate whether the step in question supports the predicted answer in context of the given CoT on a 1–5 Likert scale (Likert, 1932). We provide more details of the user study, data selection and the protocol in Appendix I.
@@ 7 Conclusion
The implementation of our proposed framework has a number of limitations, both in design as well as implementation. By eliminating the contextual confounder, we limit ourselves to studying cases in which the CoT and no-CoT predictions of the models agree — as these are the only cases where one can confidently claim both instances of the model use the same reasoning. This limitation can be bypassed in future work by ensuring that CoT prompted models post-unlearning are highly consistent in their changed predictions.
Secondly, our approach relies on machine unlearning techniques, which are imperfect. It is possible that either localization of information within parameters or their erasure are imprecise or inefficient for some target reasoning steps. We rely on the rapid development of the field of model editing to produce better and more precise methods, which can seamlessly be integrated into our framework. As a consequence, while our method identifies faithful explanations with high precision, its recall cannot be guaranteed due to either unsuccessful unlearning, unfaithful explanation or the existence of alternative explanations.
Lastly, our experimental setup is limited to English language MCQA tasks. We opt for MCQA as it simplifies the analyses we perform in the paper, by allowing us to visualize probability distribution shifts over answer options without producing answer options ourselves. Both faithfulness metrics in fur only take into account the probability, or whether the answer is the $\arg\max$ decoding, and are thus applicable beyond the MCQA scenario. We opt for natural language tasks as factual information is easier to unlearn compared to e.g. arithmetic reasoning.
@@ Appendix A Background on Machine Unlearning
Motivated by the need to erase sensitive information from machine learning models (Cao and Yang, 2015; Harding et al., 2019; Ippolito et al., 2023), the field of machine unlearning emerged as an efficient alternative to filter-then-retrain-based approaches (Neel et al., 2021; Jang et al., 2023; Eldan and Russinovich, 2023; Liu et al., 2024, inter alia). Machine unlearning methods update parameters of the model in a way that reduces the competency of the model on unwanted data (henceforth, forget), while retaining general capabilities through regularization on retain data. Such methods decrease token probabilities on the forget data by gradient-based approaches (Jang et al., 2023; Eldan and Russinovich, 2023; Gandikota et al., 2024; Li et al., 2024; Zhang et al., 2024b) or directly updating parameters (Meng et al., 2022, 2023; Wu et al., 2023; Ashuach et al., 2024).
@@
$\mathcal{L}_{\text{NPO},\beta}(\theta)=\frac{2}{\beta}\mathbb{E}_{\mathcal{D}_{\text{FG}}}\left[\text{log}\left(1+\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\beta}\right)\right],$ (7)
@@ Appendix B Alternatives to Measuring (Un)Faithfulness
A number of methods have been proposed with the goal of measuring faithfulness of model reasoning (Lanham et al., 2023; Atanasova et al., 2023; Bentham et al., 2024; Siegel et al., 2024; Chen et al., 2024; Madsen et al., 2024, inter alia). However, not all of them are applicable to our setup. Some of the methods are designed for NLI tasks (Atanasova et al., 2023; Siegel et al., 2024; Parcalabescu and Frank, 2024) while others aim at multi-hop QA tasks (Parcalabescu and Frank, 2024; Lanham et al., 2023; Chen et al., 2024; Bentham et al., 2024). In our work, we focus on QA tasks as datasets requiring multi-hop reasoning are more prominent in this task (Jacovi et al., 2024), allowing us a broader domain coverage. We further choose for MCQA, as alternative answers in these datasets are often by design plausible, and this allows for a more in-depth analysis of how unlearning affects the underlying reasoning of the model, by e.g. making it opt for plausible alternatives. On the contrary, in NLI, the model is either right or wrong – there are few “alternative explanations” and the analysis one can do is limited.
@@
We have considered other baselines applicable to CoT reasoning (cf. Table 1. in Zaman and Srivastava, 2025). Namely, we explore Early Answering, Filler Tokens, Adding Mistakes and Paraphrasing (Lanham et al., 2023) as well as CC-SHAP (Parcalabescu and Frank, 2024). Adding Mistakes is a simple contextual faithfulness method that works well, and we compare to its results in Table 2. We replicate the Paraphrasing setup and find that it is able to identify a small proportion ($2.84$%, on average) of instances as unfaithful. See full results in Table 4. Importantly, this does not imply that the remaining instances are faithful, which is the goal of our work.
@@
The remaining methods from Lanham et al. (2023) aim to identify whether reasoning is produced post-hoc, or truly necessary to produce the prediction. Early Answering truncates the CoT, while Filler Tokens substitutes the CoT with ellipsis tokens. Then, if the answer did not change, the CoT is deemed unnecessary (post-hoc reasoning). These measures do not aim to determine faithfulness of CoTs. Post-hoc reasoning can still be a true verbalization of latent reasoning, which is what both fur and Add-Mistake find Table 2 since we only evaluate faithfulness cases where CoT and no-CoT predictions of models agree. Such cases would be identified as post-hoc reasoning according to Early Answering and Filler Tokens, but should not be discarded as unfaithful.
Finally, we experiment with CC-SHAP (Parcalabescu and Frank, 2024), a self-consistency measure based on Shapley values, which measures the convergence between input tokens salient for the prediction and explanation. We use the official implementation from the authors [^9] https://github.com/Heidelberg-NLP/CC-SHAP, but when applying SHAP to instances from our dataset, relative importances of tokens from input are frequently exactly zero (importances for reasoning do not behave in this manner), which results in NaN CC-SHAP scores in $90.24$% of instances across datasets and models. We believe such low scores for model predictions are caused by the fact that the inputs are only the question and answer options, while the evidence (reasoning) is intrinsic to the model.
@@ Appendix C Dataset and Model Statistics
We report the base performance of the analyzed models on the datasets we selected, with and without CoT in Table 5. Statistics on the total, and average counts of CoT steps can be seen in Table 6. We describe and exemplify the prompting setup in Appendix D.
@@ Appendix D MCQA Task Prompts
We use two flavors of prompts when producing model predictions and the CoT for the evaluated tasks. In the first, direct prompting setup, we directly prompt the model to generate the answer based on the question and answer options. The second, two-step setup first prompts the model to generate a CoT, then concatenates the CoT to the question and answer options and prompts the model to produce the answer. Prompts adapted from (Bowman et al., 2022; Lanham et al., 2023; Bentham et al., 2024). We conduct both prompting setups in zero-shot manner.
@@
{mdframed}
[backgroundcolor=blue!5,skipabove=0.5] Human: Question: [Question]
Choices:
[Answer_choices]
Assistant: The single, most likely answer is (
@@
{mdframed}
[backgroundcolor=blue!5,skipabove=0.5] Human: Question: [Question]
Choices:
[Answer_choices]
Assistant: Let’s think step by step:
@@
{mdframed}
[backgroundcolor=blue!5,skipabove=0.5] Human: Question: [Question]
Choices:
[Answer_choices]
[Chain_of_thought]
Human: Given all of the above, what’s the single, most likely answer?"
Assistant: The single, most likely answer is (

### v3

@@ 6.1 Effectiveness of Unlearning
We report the results of unlearning when using the best hyperparameters for each method and dataset in Table˜1. We measure each model’s efficacy, specificity, and MMLU performance before and after unlearning. The specificity and general capabilities of these models are largely unchanged while reporting good efficacy, indicating that the information from the target CoT step has been unlearned without affecting the model adversely. We report the results of various learning rates and discuss methodological choices in Appendix˜E.
@@ 6.2 Does Unlearning Change Predictions?
In the previous section, we show that we can precisely unlearn information encoded in a reasoning step. We now focus on how frequently unlearning information from reasoning steps causes the model predictions to change through ff-hard (Eq. 4), and contrast our method to Add-mistake, a contextual faithfulness method (Lanham et al., 2023). Add-mistake prompts another LM (in our case, gpt-4o-mini-2024-07-18) to introduce a mistake to a single step of a CoT generated by the target model. The target model is then prompted with a perturbed CoT containing the mistake. If the prediction of the model changes, the CoT is considered faithful. We report the results of instance-level faithfulness for fur and Add-mistake in Table˜2.888We explore and comment other baselines in Appendix B. We compare to this popular baseline, as ground-truth faithfulness is not available and cannot be directly labeled since the model’s internal process is unknown. Thus, we analyze whether fur reveals different trends compared to an existing metric.
@@
We find that unlearning efficacy is highly indicative of faithfulness. The Pearson correlation between average efficacy and ff-hard is high: $0.889$ with $p<0.0001$. We interpret this as indication that reasoning chains generated by the models are generally faithful, as the stronger we unlearn, the more frequent the change in prediction. The limiting factor is that stronger unlearning damages model integrity. Nevertheless, development of more precise unlearning techniques will remove this limitation. We discuss this further, along with step-level faithfulness Appendix˜I.
@@ 6.3 Does Unlearning Change Reasoning?
Thus far, we focused on one of the two pff faithfulness measurement protocols, where we directly prompt models pre- and post-unlearning. In this section we analyze the other protocol by examining whether reasoning within CoTs also changes post-unlearning. To illustrate this, Figure˜3 visualizes how prediction probabilities of the no-CoT-prompted model change through unlearning iterations, along with the CoTs of the unlearned model. ‘Base’ refers to the model pre-unlearning. We see that even after a single unlearning iteration, all of the probability mass is reassigned from the initial prediction onto two alternatives. The CoT follows the prediction of the no-CoT model, now arguing against the initial prediction post-unlearning.
To quantitatively assess how frequently the verbalized reasoning of the model changes post-unlearning, we employ an LLM-as-a-judge (Zheng et al., 2023a) to verify if unlearning caused the generated CoT to support a different answer, indicating deeper unlearning, or if the change in model prediction is not reflected in reasoning Cohen et al. (2024). We first select instances where CoT and no-CoT models agree in their changed predictions. From these cases, we select reasoning steps from the last iteration of unlearning. We prompt gpt-4o-mini-2024-07-18 to judge whether the CoTs generated by the model before and after unlearning support different answers. We report results in Table˜3 and detail our setup in Appendix˜H.
Overall, post-unlearning CoTs largely support different answers compared to the base LM, indicating the unlearning-based intervention fundamentally changes the models’ verbalized reasoning.
Figure 4: Heatmap produced by unlearning reasoning steps. $\Delta p$ denotes $f$-soft: the change in initial answer probability. Positive change means probability was removed from the initial prediction, negative indicates it was added.
@@ 6.4 Quantifying Step Level Faithfulness
In this section, we showcase how ff-soft (Eq. 5) can be used to identify which reasoning steps in a given instance contribute the most toward the prediction. For one example in Figure˜4 we plot heatmaps for each reasoning step, which indicate how much probability mass has been shifted to (red) or from (green) the models’ initial answer when that step was unlearned. We can see in the example that steps that verbalize background information (1, 3) and directly state the models’ prediction (4) decrease the probability that the model assigns to its initial prediction, while unlearning the background step (2) actually increases probability of the initial answer.
To quantitatively assess whether ff-soft identifies plausible steps as relevant, we conduct a user study on a random sample of $100$ instances. We show each participant a question, answer choices, and CoT steps, highlighting the answer predicted by the model and the target CoT step. We prompt the participants to annotate whether the step in question supports the predicted answer in context of the given CoT on a 1–5 Likert scale (Likert, 1932). We provide more details of the user study, data selection and the protocol in Appendix˜J.
@@
7 Discussion and Future Outlook
Recent works have turned to exploring test-time scaling via reasoning language models such as OpenAI o1 (Jaech et al., 2024) and DeepSeek R1 (DeepSeek-AI et al., 2025). Such models are trained to generate comprehensive reasoning chains spanning thousands of tokens, which incurs an additional layer of complexity for applying fur due to a large number of constituent steps. While a body of work strives to reduce overthinking in CoTs (Jin et al., 2024; Hassid et al., 2025; Xu et al., 2025; Amiri et al., 2025), the problem of length still persists in cases where long CoTs are necessary. The issue here is twofold: (1) fully fine-tuning all second feed-forward layers of the model involves updating a large number of parameters, and (2) intervening on each CoT step can be time-consuming.
Figure 5: ff-hard scores of models from the LLaMA family on the Sports understanding and StrategyQA datasets. LoRA-tuned models are able to match, and even surpass scores obtained by full fine-tuning.
Reducing time complexity.
fur can innately be applied to individual CoT steps in parallel. However, not all CoT steps are equally salient for the prediction. We envision that models similar to verifiers (Lightman et al., 2024; Chen et al., 2024a; Jacovi et al., 2024) can detect, and rank, CoT steps most important for the models’ reasoning (thought anchors; Bogdan et al., 2025), which can then be prioritized for erasure. Recent work supporting our vision delves into analyzing the importance of, as well as controlling LMs chains-of-thought (Lee et al., 2025; Xiao et al., 2025; Yu et al., 2025). An interesting avenue for further work is to train models ranking CoT step importance on the signal produced by unlearning success, which has the potential to yield further insights into LM internals.
Reducing space complexity.
To address the issue of space, we explore using LoRA (Hu et al., 2022) to reduce the number of active parameters. We apply LoRA tuning with ranks $8$, $32$ and $128$ to the FF2 matrix of two LLaMA models (3B and 8B) and $30$ instances from Sports and StrategyQA. As seen in Figure˜5, LoRA offers a potential alternative to full fine-tuning. Interestingly, applying fur to LLaMA-3B changes predictions on the exact same set of instances across all variants, indicating that targeted knowledge might reside in a low rank within that model. We detail the experimental setup and provide additional results in Appendix˜F.
@@ 8 Conclusion
The implementation of our proposed framework has a number of limitations, both in design as well as implementation. By eliminating the contextual confounder, we limit ourselves to studying cases in which the CoT and no-CoT predictions of the models agree — as these are the only cases where one can hypothesize both instances of the model use the same reasoning. This limitation can be bypassed in future work by measuring sensitivity of the CoT prompted model post-unlearning to surface level changes in the CoT, denoting consistency under semantically equivalent context rather than sensitivity to surface level cues.
Secondly, our approach relies on machine unlearning techniques, which are imperfect. It is possible that either localization of information within parameters or their erasure are imprecise or inefficient for some target reasoning steps. We rely on the rapid development of the field of machine unlearning and model editing to produce better and more precise methods such as CRISP Ashuach et al. (2025a) and PISCES (Gur-Arieh et al., 2025), which can seamlessly be integrated into our framework. While our method identifies faithful explanations with high precision, its recall cannot be guaranteed due to either unsuccessful unlearning, unfaithful explanation or the existence of alternative explanations. Furthermore, applying machine unlearning requires the capability to fine-tune the target model, which makes FUR not applicable to closed API-based models. Despite this limitation, we strongly believe that credible faithfulness of natural language explanations such as CoTs requires parameter access and interventions.
Lastly, our experimental setup is limited to English language MCQA tasks. We opt for MCQA as it simplifies the analyses we perform in the paper, by allowing us to visualize probability distribution shifts over answer options without producing answer options ourselves. Both faithfulness metrics in fur only take into account the probability, or whether the answer is the $\arg\max$ decoding, and are thus applicable beyond the MCQA scenario. Applying our method to other tasks such as long-form generation can be done by assessing whether the direct answer changes after unlearning a reasoning step. We opt for natural language tasks as factual information is conceptually easier to unlearn compared to, e.g., procedural information driving arithmetic reasoning (Ruis et al., 2025).
@@ Appendix A Background on Machine Unlearning
Motivated by the need to erase sensitive information from machine learning models (Cao and Yang, 2015; Harding et al., 2019; Ippolito et al., 2023), the field of machine unlearning emerged as an efficient alternative to filter-then-retrain-based approaches (Neel et al., 2021; Jang et al., 2023; Eldan and Russinovich, 2023; Liu et al., 2024, inter alia). Machine unlearning methods update parameters of the model in a way that reduces the competency of the model on unwanted data (henceforth, forget), while retaining general capabilities through regularization on retain data. Such methods decrease token probabilities on the forget data by gradient-based approaches (Jang et al., 2023; Eldan and Russinovich, 2023; Gandikota et al., 2024; Li et al., 2024; Zhang et al., 2024b) or directly updating parameters (Meng et al., 2022, 2023; Wu et al., 2023; Ashuach et al., 2025b).
@@
$\mathcal{L}_{\text{NPO},\beta}$ $(\theta)=$ (7)
$\frac{2}{\beta}\mathbb{E}_{\mathcal{D}_{\text{FG}}}\left[\text{log}\left(1+\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\beta}\right)\right],$
@@ Appendix B Alternatives to Measuring (Un)Faithfulness
A number of methods have been proposed with the goal of measuring faithfulness of model reasoning (Lanham et al., 2023; Atanasova et al., 2023; Bentham et al., 2024; Siegel et al., 2024; Chen et al., 2024b; Madsen et al., 2024, inter alia). However, not all of them are applicable to our setup. Some of the methods are designed for NLI tasks (Atanasova et al., 2023; Siegel et al., 2024; Parcalabescu and Frank, 2024) while others aim at multi-hop QA tasks (Parcalabescu and Frank, 2024; Lanham et al., 2023; Chen et al., 2024b; Bentham et al., 2024). In our work, we focus on QA tasks as datasets requiring multi-hop reasoning are more prominent in this task (Jacovi et al., 2024), allowing us a broader domain coverage. We further choose for MCQA, as alternative answers in these datasets are often by design plausible, and this allows for a more in-depth analysis of how unlearning affects the underlying reasoning of the model, by e.g. making it opt for plausible alternatives. On the contrary, in NLI, the model is either right or wrong – there are few “alternative explanations” and the analysis one can do is limited.
@@
We have considered other baselines applicable to CoT reasoning (cf. Table 1. in Zaman and Srivastava, 2025). Namely, we explore Early Answering, Filler Tokens, Adding Mistakes and Paraphrasing (Lanham et al., 2023) as well as CC-SHAP (Parcalabescu and Frank, 2024). Adding Mistakes is a simple contextual faithfulness method that works well, and we compare to its results in Table˜2. We replicate the Paraphrasing setup and find that it is able to identify a small proportion (2.842.84%, on average) of instances as unfaithful. See full results in Table˜4. Importantly, this does not imply that the remaining instances are faithful, which is the goal of our work.
@@
The remaining methods from Lanham et al. (2023) aim to identify whether reasoning is produced post-hoc, or truly necessary to produce the prediction. Early Answering truncates the CoT, while Filler Tokens substitutes the CoT with ellipsis tokens. Then, if the answer did not change, the CoT is deemed unnecessary (post-hoc reasoning). These measures do not aim to determine faithfulness of CoTs. Post-hoc reasoning can still be a true verbalization of latent reasoning, which is what both fur and Add-Mistake find Table˜2 since we only evaluate faithfulness cases where CoT and no-CoT predictions of models agree. Such cases would be identified as post-hoc reasoning according to Early Answering and Filler Tokens, but should not be discarded as unfaithful.
Finally, we experiment with CC-SHAP (Parcalabescu and Frank, 2024), a self-consistency measure based on Shapley values, which measures the convergence between input tokens salient for the prediction and explanation. We use the official implementation from the authors [^9] https://github.com/Heidelberg-NLP/CC-SHAP, but when applying SHAP to instances from our dataset, relative importances of tokens from input are frequently exactly zero (importances for reasoning do not behave in this manner), which results in NaN CC-SHAP scores in 90.2490.24% of instances across datasets and models. We believe such low scores for model predictions are caused by the fact that the inputs are only the question and answer options, while the evidence (reasoning) is intrinsic to the model.
@@ Appendix C Dataset and Model Statistics
We report the base performance of the analyzed models on the datasets we selected, with and without CoT in Table˜5. Statistics on the total, and average counts of CoT steps can be seen in Table˜6. We describe and exemplify the prompting setup in Appendix˜D.
@@ Appendix D MCQA Task Prompts
We use two flavors of prompts when producing model predictions and the CoT for the evaluated tasks. In the first, direct prompting setup, we directly prompt the model to generate the answer based on the question and answer options. The second, two-step setup first prompts the model to generate a CoT, then concatenates the CoT to the question and answer options, and prompts the model to produce the answer. Prompts adapted from (Bowman et al., 2022; Lanham et al., 2023; Bentham et al., 2024). We conduct both prompting setups in zero-shot manner.
@@
Human: Question: [Question] Choices: [Answer_choices] Assistant: The single, most likely answer is (
@@
Human: Question: [Question] Choices: [Answer_choices] Assistant: Let’s think step by step:
@@
Human: Question: [Question] Choices: [Answer_choices] [Chain_of_thought] Human: Given all of the above, what’s the single, most likely answer?" Assistant: The single, most likely answer is (